Shift Summary Sept 21, 1999 Day

Douglas A. Jensen (djensen@fnal.gov)
Tue, 21 Sep 1999 15:48:55 -0500 (CDT)

Shift Summary for DAY shift, Sept 21, 1999

On Shift: Coleman, Jensen
and assorted visitors, Hogan, Jason (L&H), BobT, Sydney,

Hours of beam available at all intensities: 5.5 /8.5
Hours of beam available at nominal intensity: 5.5
Hours of E799 data (special runs not included): 4.0
Hours of special runs (laser scans, muon runs...): 1.5

Accounting of lost time:
No beam first 2 hours, No beam 45 min for NuMI access, missed spills etc
Worked on hyperon trigger, and getting to 1E13

========================================================================
Runs ended this shift:

14655, 14656
========================================================================
Tapes brought to FCC this shift:
(including tapes for online split)
UPA007-UPA015, UPB007-UPB015, UPC007-UPC015, UPS003-UPS005

========================================================================
Accesses during shift:None

========================================================================
Overview of shift:
Run 14655 in progress, but there has been no beam for hours at start.
Continued taking data, starting run 14656 still at 8E12. Off for
45 min for Numi access, then go to 1E13 and do some hyperon trigger
studdies ( kaon triggers stable )
At end of shift - start clean data taking run, hyperons=>hyperons/2
for the time being.
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14655 07:00 New Shift arrives to wait for beam:
LRF problems, no estimate (6:56 message )
'95' 08:30 Call MCR, they est 45 - 60 min to beam.
117 09:00 Start getting beam, they offer 0.5 nominal
intensity for now.
130 09:17 Getting 8E12 now
243 11:48 KFRIENDS - latches lost; pipeline,.. 2-3 spills lost
256 12:03 triggers stop

Mounted UPF002 on tps3d2 KTeV2 as requested

264 12:15 'Tapes full', run ends

14656 1 12:27 start run, with pipeline err,reset with DBC
KFRIEND, resets....
6 12:33 getting data. 8E12
29 13:03 Beam off for kicker controls change for NuMI
target tests , est 30 min
59 13:45 Beam back.
73 14:00 End run,

new intensity request of 1.0E13 is being implemented

14657 5 14:15 Running, after 3 kfriends... Intens Req 1E13
will be a garbage run, as tweeking hyperon trigger
'on the fly', by Arun, under supervision of BobT
Add hc2xloose to hyperon
15 14:30 running again. checking scaler rates....
did not lower rate hardly at all.
29 Add HC2Xtight - lowers rate ( but dangerous!? )

55 +/- Go to a nominal ( as we have been running ) hyperon
trigger - prescaled by 2. HCC>=2 in hyperon trig.
no HAveto, 2HCYloose, no X hit counting.
Assorted aborted spills. Tricky to understand
55 15:25 End run ( time to do physics )

14658 3 15:40 Running


Note - spill number in ' ' implies 'virtual' spill; i.e. no beam.
77.88 sec/spill = 1/(machine cycle time)
========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift(s)
Xerox machine still not available. Have not heard anything in a
long long long time. There is a note that it will be fixed soon, but
the note is showing clear signs of aging - becoming indistinguishable
from a piece of the 'Dead Sea Scrolls'.